OpenStudy (anonymous):

will give medals and fan What is the perimeter of this parallelogram? A parallelogram is shown with the bases each 8 inches, the height 2 inches, and the length of one of the sides 3 inches

OpenStudy (anonymous):

@joseg.77 could u help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where is your figure?

OpenStudy (anonymous):



OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

oppsosite sides of a parallelogram are congruent Perimeter of the parallelogram = 8+3+8+3= 22 inches

OpenStudy (anonymous):

What are options?

OpenStudy (anonymous):

thirteen inches sixteen inches twenty-two inches twenty-four inches

OpenStudy (anonymous):

oppsosite sides of a parallelogram are congruent Perimeter of the parallelogram = 8+3+8+3= 22 inches

OpenStudy (anonymous):

@nova2528

OpenStudy (anonymous):

thnx the answer is 22 then

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

22